# Mimicking Human Intuition: Cognitive Belief-Driven Reinforcement Learning

### v3

@@ 5 Algorithm Implementation
Algorithmic instantiations of the Cognitive Belief-Driven Reinforcement Learning (CBD-RL) framework include Cognitive Belief-Driven Q-learning (CBDQ) for value-based, and its policy-based counterparts—Cognitive Belief-Driven Soft Actor-Critic (CBDSAC) and Cognitive Belief-Driven Proximal Policy Optimization (CBDPPO).
5.1 Cognitive Belief-Driven Q-learning (CBDQ)
CBDQ operationalizes the CBD-RL framework within a discrete action domain by integrating short-term reward learning with long-term conceptual preferences. Rather than relying on greedy action selection, CBDQ applies a smoothed Bellman operator where the next-state value estimate is computed using a belief-preference distribution. This distribution merges reward-informed action likelihoods with concept-level historical priors derived from Conceptual Category Formation (CCF).
For a given state $s_{t}$, its conceptual category $C_{k}$ is identified via distance-based clustering. A category-specific belief distribution $\mathcal{P}_{k}(a\mid s_{t})$ is constructed using normalized action frequencies. This prior is fused with $q_{t}(a\mid s_{t+1})$ to form:
$b_{t}(a\mid s_{t+1})=(1-\beta_{t})\cdot q_{t}(a\mid s_{t+1})+\beta_{t}\cdot\mathcal{P}_{k}(a\mid s_{t}),$ (12)
where $\beta_{t}\in[0,1]$ is an adaptive parameter controlling the influence of conceptual memory.
The Q-function update rule becomes:
$\begin{split}Q_{t+1}(s_{t},a_{t})=\ Q_{t}(s_{t},a_{t})+\alpha_{t}(s_{t},a_{t})\Big{[}r(s_{t},a_{t})\\ +\gamma\sum_{a\in A}b_{t}(a\mid s_{t+1})Q_{t}(s_{t+1},a)-Q_{t}(s_{t},a_{t})\Big{]}\end{split}$ (13)
where the expectation over $b_{t}$ acts as a smoothed substitute for the max operator in classical Q-learning, enabling uncertainty-aware value propagation.
Practically, CBDQ uses an $\epsilon$-greedy exploration strategy with actions sampled from $b_{t}(a\mid s_{t+1})$. The belief model is implemented as an online frequency table per category, updated incrementally with each transition. Experience replay is used to stabilize training, and conceptual categories are periodically re-evaluated to ensure semantic coherence.
This mechanism allows CBDQ to leverage both immediate task feedback and accumulated structural knowledge, promoting faster convergence and improved generalization. The full procedure is described in Appendix A.1, and convergence is discussed in Appendix C under the Smoothed Bellman formulation.
5.2 Cognitive Belief-Driven Soft Actor-Critic (CBDSAC)
To support concept-informed decision-making in continuous action spaces, we integrate the Cognitive Belief-Driven Reinforcement Learning (CBD-RL) framework into Soft Actor-Critic (SAC), forming the Cognitive Belief-Driven Soft Actor-Critic (CBDSAC). While SAC optimizes a stochastic Gaussian policy through entropy-regularized objectives, it relies solely on immediate feedback, lacking mechanisms for retaining and transferring structured knowledge across semantically similar but temporally distant contexts.
CBDSAC augments SAC by incorporating belief distributions derived from Conceptual Category Formation (CCF), which cluster states into semantically coherent groups and provide structured priors over actions. Instead of relying solely on the actor network’s output, CBDSAC fuses its learned policy distribution with a corresponding conceptual belief, previously constructed and updated using Bayesian statistics (see Chapter 4). This fusion yields a belief-preference distribution that modulates both action sampling and policy updates. The actor network outputs a Gaussian policy:
$\mathcal{Z}_{t}(\cdot\mid s_{t+1})=\mathcal{N}(\mu_{\pi_{\theta}}(s_{t+1}),\sigma^{2}_{\pi_{\theta}}(s_{t+1})),$ (14)
which captures short-term, reward-driven action preferences. CBDSAC fuses this distribution with the conceptual belief to construct a belief-preference distribution:
$\mu_{\text{blend}}(s_{t+1})$ $=(1-\beta_{t})\mu_{\pi_{\theta}}(s_{t+1})+\beta_{t}\mu_{\text{post}}(s_{t}),$ (15)
$\sigma^{2}_{\text{blend}}(s_{t+1})$ $=(1-\beta_{t})\sigma^{2}_{\pi_{\theta}}(s_{t+1})+\beta_{t}\sigma^{2}_{\text{post}}(s_{t}),$ (16)
where $\beta_{t}\in[0,1]$ modulates the influence of structured belief versus reactive policy.
This yields the final action distribution:
$b_{t}(\cdot\mid s_{t+1})=\mathcal{N}(\mu_{\text{blend}}(s_{t+1}),\sigma^{2}_{\text{blend}}(s_{t+1})),$ (17)
which governs both action sampling and gradient-based updates.
Policy optimization proceeds via entropy-regularized soft policy iteration:
$\mathbb{E}_{s\sim\mathcal{D},a\sim b_{t}}\left[Q_{\phi}(s,a)-\alpha\log b_{t}(a\mid s_{t+1})\right],$ (18)
with:
$\log b_{t}(a\mid s_{t+1})=-\frac{1}{2}\Bigg{(}\frac{(a-\mu_{\text{blend}})^{2}}{\sigma_{\text{blend}}^{2}}+2\log\sigma_{\text{blend}}+c\Bigg{)}.$ (19)
where $c$ is a fixed constant.
By integrating policy learning with semantically grounded beliefs, CBDSAC enables agents to generalize across conceptually coherent behaviors. This fusion facilitates better sample reuse, long-term coherence, and more human-like decision-making. Full implementation details and pseudocode are provided in Appendix A.2.
5.3 Cognitive Belief-Driven Proximal Policy Optimization (CBDPPO)
To incorporate concept-level priors into on-policy learning, the Cognitive Belief-Driven Reinforcement Learning (CBD-RL) framework is integrated with the clipped surrogate formulation of Proximal Policy Optimization (PPO), yielding the Cognitive Belief-Driven PPO (CBDPPO) algorithm. While PPO maximizes an advantage-weighted likelihood ratio under a trust region constraint (Qiao et al., 2023), the policy update remains driven solely by immediate feedback, limiting its ability to leverage structural regularities observed across semantically related trajectories.
CBDPPO addresses this limitation by blending the current policy $\pi_{\theta}(a\mid s)$ with the conceptual belief $P_{k}(a\mid s)$ associated with the category $C_{k}$ that contains state $s$. The resulting belief-preference policy is defined as:
$b_{t}(a\mid s)=(1-\beta_{t})\pi_{\theta}(a\mid s)+\beta_{t}P_{k}(a\mid s),$ (20)
@@ 5.3 Cognitive Belief-Driven Proximal Policy Optimization (CBDPPO)
where the scheduling parameter $\beta_{t}\in[0,1]$ controls the influence of concept priors and increases gradually throughout training. The clipped surrogate objective of CBDPPO is:
$\mathcal{L}_{\text{CBDPPO}}=\mathbb{E}_{(s,a)\sim\pi_{\theta_{old}}}\Big{[}\min\Big{(}\frac{b_{t}(a\mid s)}{\pi_{\theta_{old}}(a\mid s)}A_{t},$ (21)
$\qquad\quad\;\text{clip}\left(\frac{b_{t}(a\mid s)}{\pi_{\theta_{old}}(a\mid s)},1-\epsilon,1+\epsilon\right)A_{t}\Big{)}\Big{]},$
@@
where $A_{t}$ is the advantage estimate and $\epsilon$ controls the trust region width. The critic and entropy terms follow the original PPO formulation; gradients are propagated through $b_{t}$, allowing concept priors to steer policy updates while the clip operator guarantees trust-region stability. The critic and entropy terms follow the original PPO formulation; gradients are propagated through $b_{t}$, allowing concept priors to steer policy updates while the clip operator guarantees trust-region stability. Full implementation details and pseudocode are provided in Appendix A.3.
@@ 6 Experiment
Running Setting
Evaluation is based on Feasible Cumulative Rewards, where higher values indicate better performance, averaged over three seeds (123, 321, 666). Conceptual clustering within CCF is simulated via clustering algorithms that group similar state-action pairs into latent categories (Appendix). All methods use identical hyperparameters and are implemented on the XuanCe benchmark suite (Liu et al., 2023), with full configurations in Appendix E.4.
Comparison Methods
For discrete action spaces, the comparison includes CBDQ and CBDPPO and the following baselines: (1) DQN (Mnih et al., 2013) with neural networks for Q-value approximation; (2) DDQN with decoupled action selection and evaluation; (3) DuelDQN with separate state value and action advantage estimation; (4) PPO with clipped objective for stable policy optimization. For continuous action spaces, CBDSAC is compared with: (1) A2C (Mnih, 2016) with synchronized advantage estimation; (2) PPO (Schulman et al., 2017) with clipped surrogate objective; (3) SAC (Haarnoja et al., 2018) with entropy-regularized policy optimization; (4) DDPG (Lillicrap, 2015) with deterministic policy gradient and actor-critic architecture.
6.1 Empirical Evaluations in Discrete Action Space
The evaluation covers a broad spectrum of environments, including Classic Control, Box2D, MetaDrive (Li et al., 2022), and Atari domains. These span from simple control tasks to complex, human-like decision-making scenarios with high-dimensional inputs and dynamic variability. CBDQ demonstrates clear advantages in structured control settings, consistently achieving faster convergence and higher final returns by leveraging belief modeling and conceptual abstraction. In more visually complex environments such as Atari, CBDPPO outperforms PPO by effectively integrating belief-guided priors into policy updates, leading to improved sample efficiency and policy robustness.
Table 1: Reward results for different algorithms on continuous action environments
Environment/Method CBDSAC SAC DDPG PPO A2C

### v4

@@ 5 Algorithm Implementation
We apply HI-RL framework into multiple RL paradigms by developing HI-Q, HI-PPO and HI-SAC (For HI-TD3, see Appendix A.1).
5.1 Conceptual Belief-Informed Q-learning (HI-Q)
The classical Deep Q-learning (DQN) algorithm (mnih2015human) relies on updating the Q-function network using the greedy Bellman operator. Namely, in any iteration $t$, with the sampled batch $D_{t}$ and any sample $(s_{i},a_{i},r_{i},s_{i}^{\prime})\in D_{t}$ within it, the learning target of the Q-network $Q_{\theta_{t+1}}(s_{i},a_{i})$ would be
$r_{i}+\gamma\max_{a\in\mathcal{A}}Q_{\theta_{t}}(s_{i}^{\prime},a_{i}),$ (4)
where $\theta_{t}$ represent the Q-network parameter at time $t$. With DQN in mind, we propose HI-Q to replace the learning target to a new one combining both the current Q-network information $Q_{\theta_{t}}$ and the conceptual abstraction experience prior $b_{t}$.
Specifically, we first introduce the construction of the concept-based belief prior $b_{t}(\cdot\,|\,k)$ at each time step $t$. Here, $b_{t}(\cdot\,|\,k)$ will be defined as the action visiting frequency summarzied over all state within the concept set $C_{k}$. For the discrete finite action space $\mathcal{A}$, we denote the number of visiting time over each state-action pair at time step t as $N_{t}(s,a)$. Then the experience prior $b_{t}$ of any $k$-th concept will be constructed as
$\forall(a,k)\in\mathcal{A}\times[K]:\quad b_{t}(a\mid k)=\frac{\sum_{s\in C_{k}}N_{t}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}\sum_{s\in C_{k}}N_{t}(s,a^{\prime})}.$ (5)
The update of $b_{t}$ is typically computational easily, since upon executing an sample tuple $(s_{i},a_{i},r_{i},s_{i}^{\prime})$, only the $(s_{i},a_{i})$-associated concept $b_{t}(a_{i}\,|\,c(s_{i}))$ will be updated.
Therefore, the combined information for any sample tuple $(s_{i},a_{i},r_{i},s_{i}^{\prime})$ associated with state $s_{i}^{\prime}$ at time $t$ is defined as
$B_{t}(\cdot\mid s_{i}^{\prime})=(1-\beta_{t})q_{t}(\cdot\mid s_{i}^{\prime})+\beta_{t}b_{t}(\cdot\,|\,c(s_{i}^{\prime})),$ (6)
where $\beta_{t}$ is a dynamic coefficient and $q_{t}(\cdot\mid s^{\prime}_{i})$ denotes the task-driven action-preference distribution, typically instantiated as a smoothing distribution over $Q$-values (e.g., softmax with temperature $\tau_{t}$ or clipped-max with exploration mass $\delta_{t}$) that gradually concentrates on the greedy action as $t$ increases (barber2023smoothed), computed via a softmax over the top-$k$ Q-values of state $s^{\prime}_{i}$, effectively assigning higher probabilities to the most promising actions:
$q_{t}(a\mid s^{\prime}_{i})=\frac{\exp(Q(s^{\prime}_{i},a)/\tau)}{\sum_{a^{\prime}\in\text{top-}k(s^{\prime}_{i})}\exp(Q(s^{\prime}_{i},a^{\prime})/\tau)},\quad a\in\text{top-}k(s^{\prime}_{i})$ (7)
where $\tau$ is softmax temperature constant. With the constructed concept-based belief based template $B_{t}$ in hand, we replace the (greedy) maximum operator in Eq. 4 of classical Q-learning to a smoothed surrogate one combining both the smoothed-greedy operator of the current Q-function and the conceptual-based belief. Therefore, the new target in HI-Q for the Q-network to learn is defined as
$r_{i}+\gamma\sum_{a\in\mathcal{A}}B_{t}(a\mid s_{i}^{\prime})Q_{t}(s_{i}^{\prime},a).$ (8)
The entire algorithm is specified in Appendix A.2.1. Our conceptual-abstraction belief enables HI-Q to leverage both immediate task feedback and accumulated conceptual structures, facilitating faster learning by borrowing experience from other similar concepts.
5.2 Conceptual Belief-Informed Proximal Policy Optimization (HI-PPO)
The standard PPO (schulman2017proximal) is a policy gradient algorithm which updates the policy by performing stochastic gradient ascent on a surrogate objective function. For any time step $t$, let $D_{t}$ be a sampled batch and $(s_{i},a_{i},r_{i},s_{i}^{\prime})\in D_{t}$ any individual sample within it. The objective is to update the policy $\pi_{\theta}(a\mid s)$ via the following loss function:
$\mathcal{L}_{\text{PPO}}=\mathbb{E}_{(s_{i},a_{i})\sim\pi_{\theta_{\text{old}}}}\Big[\min\Big(\tfrac{\pi_{\theta}(a_{i}\mid s_{i})}{\pi_{\theta_{\text{old}}}(a_{i}\mid s_{i})}A_{t},\;\text{clip}\big(\tfrac{\pi_{\theta}(a_{i}\mid s_{i})}{\pi_{\theta_{\text{old}}}(a_{i}\mid s_{i})},1-\epsilon,1+\epsilon\big)A_{t}\Big)\Big].$ (9)
where $\theta$ denotes the policy parameters, $A_{t}$ is the advantage estimate at time step $t$, and $\epsilon$ controls the trust region. While PPO updates the policy via an advantage-weighted likelihood ratio within this trust region, it depends only on immediate feedback, limiting its ability to exploit structural regularities. To overcome this, HI-PPO integrates the current policy $\pi_{\theta}(a\mid s)$ with the conceptual abstraction prior $b_{t}$.
In this paper, we focus on applying PPO in discrete action-space environments, with modifications analogous to those in HI-Q. At each time step $t$, we compute a concept-based belief prior $b_{t}(\cdot\mid k)$, defined as the action visitation frequency aggregated over all states in concept set $C_{k}$. Its computation and update follow Eq. 5, and it is combined with the policy $\pi_{\theta}(\cdot\mid s_{i})$ for state $s_{i}\in D_{t}$ at time $t$ as:
$B_{t}(\cdot\mid s_{i})=(1-\beta_{t})\pi_{\theta}(\cdot\mid s_{i})+\beta_{t}b_{t}(\cdot\mid c_{k}(s_{i})),$ (10)
@@ 5.2 Conceptual Belief-Informed Proximal Policy Optimization (HI-PPO)
where the scheduling parameter $\beta_{t}\in[0,1]$ controls the influence of concept priors and increases gradually throughout training. The clipped surrogate objective of HI-PPO is:
$\mathcal{L}_{\text{HI-PPO}}=\mathbb{E}_{(s_{i},a_{i})\sim\pi_{\theta_{\text{old}}}}\Big[\min\Big(\tfrac{B_{t}(a_{i}\mid s_{i})}{\pi_{\theta_{\text{old}}}(a_{i}\mid s_{i})}A_{t},\;\text{clip}\big(\tfrac{B_{t}(a_{i}\mid s_{i})}{\pi_{\theta_{\text{old}}}(a_{i}\mid s_{i})},1-\epsilon,1+\epsilon\big)A_{t}\Big)\Big].$ (11)
@@
The critic and entropy terms follow the original PPO formulation; gradients are propagated through $B_{t}$, allowing concept priors to steer policy updates while the clip operator guarantees trust-region stability. More implementation details and pseudocode are provided in Appendix A.2.3.
5.3 Conceptual Belief-Informed Soft Actor-Critic (HI-SAC)
Traditional Soft Actor-Critic (SAC) is a maximum entropy reinforcement learning algorithm that integrates both an actor and a critic network (haarnoja2018soft). Given a sampled batch $D_{t}$ at time step $t$, containing tuples $(s_{i},a_{i},r_{i},s_{i}^{\prime})$, the updates of the actor and critic networks parameters $\theta$ , $\phi$ from $\pi_{\theta}$ and $Q_{\phi}$ respectively are defined as follows:
$\mathcal{L}_{\text{critic}}(\phi_{i})$ $=\mathbb{E}\Big[\big(Q_{\phi_{i}}(s_{i},a_{i})-y_{t}\big)^{2}\Big],\quad i=1,2,$ (12)
where $\quad y_{t}=r_{i}+\gamma\,\mathbb{E}_{a_{i}^{\prime}\sim\pi_{\theta}}\big[Q_{\min}(s_{i}^{\prime},a_{i}^{\prime})-\alpha\log\pi_{\theta}(a_{i}^{\prime}\mid s_{i}^{\prime})\big],$
$\mathcal{L}_{\text{actor}}(\theta)$ $=\mathbb{E}_{s_{i}\sim\mathcal{D}_{t},\,a_{i}\sim\pi_{\theta}}\big[\alpha\log\pi_{\theta}(a_{i}\mid s_{i})-\min\{Q_{\phi_{1}}(s_{i},a_{i}),\,Q_{\phi_{2}}(s_{i},a_{i})\}\big].$
where $\alpha$ is entropy temperature coefficient, $y_{t}$ is the TD target, computed by the next state $s_{i}^{\prime}$ at time step $t+1$ and the corresponding action $a_{i}^{\prime}$ sampled from the policy $\pi_{\theta}$. In SAC, the Q value is computed as the minimum of the estimates from two critic networks $Q_{\phi_{i}}$ and the actor network produces a Gaussian policy in this paper:
$\pi_{\theta}(\cdot\mid s)=\mathcal{N}(\mu_{\pi_{\theta}}(s),\sigma^{2}_{\pi_{\theta}}(s)),$ (13)
where $\mu_{\pi_{\theta}}(s)$ and $\sigma^{2}_{\pi_{\theta}}(s)$ denote the mean and variance predicted by the policy network for state $s$. To support concept-informed decision-making in continuous action spaces, we propose HIS-AC to integrate current actor network $\pi_{\theta_{t}}$ and the conceptual experience prior $b_{t}$. Unlike HI-Q and HI-PPO, the concept-based belief prior $b_{t}(k)$ constructed at each time step t is defined over the actor network parameters corresponding to the states s within the concept set $C_{k}$:
$\forall(\mu,\sigma^{2},k)\in\{\mu,\sigma^{2}\}\times[K]:\quad b_{t}(k)=\{\mu_{\pi_{\theta}}(s),\sigma^{2}_{\pi_{\theta}}(s)\},\quad s\in C_{k}$ (14)
During training, we update $b_{t}$ using a Bayesian posterior update. Let $\mu_{c}$ and $\sigma_{c}^{2}$ be the parameters of the current policy $\pi_{\theta_{t}}(s)$ and $\mu_{e}$ and $\sigma_{e}^{2}$ be the experience stored in $b_{t-1}(k)$:
$b_{t}(k)=\{\frac{\sigma_{c}^{2}\,\mu_{e}+\sigma_{e}^{2}\,\mu_{c}}{\sigma_{c}^{2}+\sigma_{e}^{2}},\frac{1}{\frac{1}{\sigma_{c}^{2}}+\frac{1}{\sigma_{e}^{2}}}\},\quad(\mu_{c},\sigma_{c}^{2})\sim\pi_{\theta_{t}}(s),\quad(\mu_{e},\sigma_{e}^{2})\sim b_{t-1}(k),\quad s\in C_{k}$ (15)
At the same time, for any sample tuple $(s_{i},a_{i},r_{i},s_{i}^{\prime})$ at time step $t$, we use both $s_{i}$ and $s_{i}^{\prime}$ to obtain the corresponding $(\mu_{\pi_{\theta_{t}}}(s_{i}),\sigma^{2}_{\pi_{\theta_{t}}}(s_{i}))$ and $(\mu_{\pi_{\theta_{t}}}(s_{i}^{\prime}),\sigma^{2}_{\pi_{\theta_{t}}}(s_{i}^{\prime}))$ for the actor and critic, respectively. This fusion method can then be formally defined as:
$\mu_{\text{actor}}(s_{i})$ $=(1-\beta_{t})\mu_{\pi_{\theta_{t}}}(s_{i})+\beta_{t}\mu_{b},\quad\sigma^{2}_{\text{actor}}(s_{i})=(1-\beta_{t})\sigma^{2}_{\pi_{\theta_{t}}}(s_{i})+\beta_{t}\sigma^{2}_{b},$ (16)
$\mu_{\text{critic}}(s_{i}^{\prime})$ $=(1-\beta_{t})\mu_{\pi_{\theta_{t}}}(s_{i}^{\prime})+\beta_{t}\mu_{b},\quad\sigma^{2}_{\text{critic}}(s_{i}^{\prime})=(1-\beta_{t})\sigma^{2}_{\pi_{\theta_{t}}}(s_{i}^{\prime})+\beta_{t}\sigma^{2}_{b},$
where $(\mu_{b},\sigma^{2}_{b})\sim b_{t}(k),\quad s_{i},s_{i}^{\prime}\in C_{k}$
where $\beta_{t}\in[0,1]$ adaptively controls the relative weighting between task-driven and concept-informed signals, and $\mu_{b},\sigma^{2}_{b}$ denote the currently stored conceptual experience in $b_{t}(k)$. This results in the conceptual belief-informed distribution for both the actor and critic:
$B_{t}(\cdot\mid s_{i}=\mathcal{N}(\mu_{\text{actor}}(s_{i}),\sigma^{2}_{\text{actor}}(s_{i})),\quad B_{t}(\cdot\mid s_{i}^{\prime})=\mathcal{N}(\mu_{\text{critic}}(s_{i}^{\prime}),\sigma^{2}_{\text{critic}}(s_{i}^{\prime})).$ (17)
Finally, we replace the policy $\pi_{\theta}$ in Eq.12 with the integrated $B_{t}$ and perform the updates accordingly:
$\mathcal{L}_{\text{HI-SAC}_{critic}}(\phi_{i})$ $=\mathbb{E}\Big[\big(Q_{\phi_{i}}(s_{i},a_{i})-y_{t}\big)^{2}\Big],\quad i=1,2,$ (18)
$\text{where}\quad y_{t}$ $=r_{i}+\gamma\,\mathbb{E}_{a_{i}^{\prime}\sim B_{t}}\big[Q_{\min}(s_{i}^{\prime},a_{i}^{\prime})-\alpha\log B_{t}(a_{i}^{\prime}\mid s_{i}^{\prime})\big],$
$\mathcal{L}_{\text{HI-SAC}_{actor}}(\theta)$ $=\mathbb{E}_{s_{i}\sim\mathcal{D}_{t},\,a_{i}\sim B_{t}}\big[\alpha\log B_{t}(a_{i}\mid s_{i})-\min\{Q_{\phi_{1}}(s_{i},a_{i}),\,Q_{\phi_{2}}(s_{i},a_{i})\}\big].$
By integrating policy learning with semantically grounded beliefs, HI-SAC enables agents to generalize across conceptually coherent behaviors. This fusion facilitates better sample reuse, long-term coherence, and more human-like decision-making. The pseudocodes are provided in Appendix A.2.2.
@@ 6 Experiment
Experimental setup: Evaluation is based on Feasible Cumulative Rewards, where higher values indicate better performance, averaged over three seeds (123, 321, 666). The evaluation spans a wide range of environments, including Classic Control, Box2D (catto2005iterative), MetaDrive (li2022metadrive), MuJoCo (todorov2012mujoco), and Atari (bellemare2013arcade) domains. Conceptual clustering is simulated using clustering algorithms that group similar state-action pairs into latent categories. All methods employ identical hyperparameters and are implemented on the XuanCe benchmark suite (liu2023xuance).
Evaluated methods: For discrete action spaces, we compare HI-Q and HI-PPO with the following baselines: DQN (mnih2013playing), DDQN (van2016deep), DuelDQN (wang2016dueling), and PPO(schulman2017proximal), covering standard Q-value approximations, decoupled action evaluation, state-action advantage estimation, and clipped policy optimization. For continuous action spaces, HI-SAC is compared with A2C (mnih2016asynchronous), PPO, SAC (haarnoja2018soft), and DDPG (lillicrap2015continuous), representing common policy-gradient and actor-critic methods with entropy regularization or deterministic gradients.
Table 1: Average cumulative rewards of HI-RL variants and baselines across discrete and continuous action environments.
HI-RL for DQN Variants
Environment/Method HI-DQN PPO DQN Duel_DQN DDQN
Classic Control - CartPole 499.78 $\pm$ 0.22 499.17 $\pm$ 0.83 478.44 $\pm$ 21.56 440.69 $\pm$ 59.31 396.51 $\pm$ 103.49
Classic Control - Acrobot -80.57 $\pm$ 17.48 -500.00 $\pm$ 0.00 -87.19 $\pm$ 18.55 -104.53 $\pm$ 54.19 -100.77 $\pm$ 24.79
Box2d - CarRacing 854.66 $\pm$ 45.35 189.05 $\pm$ 56.48 830.78 $\pm$ 51.61 -13.05 $\pm$ 24.66 766.16 $\pm$ 88.22
Box2d - LunarLander 232.73 $\pm$ 40.20 204.95 $\pm$ 48.77 52.67 $\pm$ 192.08 -58.97 $\pm$ 4.08 191.79 $\pm$ 69.16
MetaDrive - rXTSC 189.22 $\pm$ 63.71 156.74 $\pm$ 31.44 82.05 $\pm$ 82.84 39.50 $\pm$ 7.27 185.55 $\pm$ 107.80
MetaDrive - TOrSX 159.39 $\pm$ 38.40 149.97 $\pm$ 26.28 101.60 $\pm$ 13.72 69.16 $\pm$ 14.07 83.77 $\pm$ 22.37
MetaDrive - XTOC 303.15 $\pm$ 50.89 293.72 $\pm$ 66.42 170.73 $\pm$ 31.60 67.42 $\pm$ 6.29 170.73 $\pm$ 31.60
MetaDrive - XTSC 233.91 $\pm$ 64.92 191.50 $\pm$ 39.31 215.94 $\pm$ 205.74 63.47 $\pm$ 4.96 147.71 $\pm$ 92.55
MetaDrive - CYrXT 97.99 $\pm$ 25.43 97.83 $\pm$ -38.66 77.23 $\pm$ 47.94 9.12 $\pm$ 39.53 75.39 $\pm$ 49.99
MetaDrive - COrXSrT 117.90 $\pm$ 24.56 89.27 $\pm$ 23.52 117.18 $\pm$ 30.28 53.01 $\pm$ 4.91 29.15 $\pm$ 16.26
MetaDrive - SrOYCtryS 130.27 $\pm$ 117.07 75.38 $\pm$ 8.12 105.01 $\pm$ 88.37 38.90 $\pm$ 0.39 100.72 $\pm$ 81.92
HI-RL for SAC Variants
Environment/Method HI-SAC SAC PPO DDPG A2C